OpenStudy (anonymous):

A cylinder labeled X has a height of 12 inches and a radius of 10 inches. Another cylinder labeled Y has a height of 36 inches and a radius of 30 inches. The volume of cylinder Y is _______ times the volume of cylinder X.

OpenStudy (anonymous):

I'm thinking 9????

jimthompson5910 (jim_thompson5910):

Volume of cylinder X V = pi*r^2*h V = pi*10^2*12 V = 1200pi ------------------- Volume of cylinder Y V = pi*r^2*h V = pi*30^2*36 V = 32400pi ------------------ Ratio of volume of cylinder Y to volume of cylinder X = 32400pi/1200pi = 27 So The volume of cylinder Y is ___27____ times the volume of cylinder X.

jimthompson5910 (jim_thompson5910):

As a shortcut, you can notice that each dimension in cylinder Y is 3 times the corresponding dimension in cylinder X. So the volume in cylinder Y will be 3^3 = 27 times greater than cylinder X.

OpenStudy (anonymous):

Thanks, I had no idea what to do :) now it's easier to understand

jimthompson5910 (jim_thompson5910):

that's great :)